# Exploring the Impact of Table-to-Text Methods on Augmenting LLM-based Question Answering with Domain Hybrid Data

### 20 Feb 2024 v1

@@ 1 Introduction
In terms of implementation, we collect a real-world hybrid dataset called ICT-DATA, by extracting text and tables from numerous documents about Information and Communication Technology (ICT) products. It is important to note that the text contained in tables accounts for approximately 18% of the total content in ICT-DATA. We employ different table-to-text methods to process the tables in ICT-DATA, obtaining different ICT corpora. These corpora are then utilized to build QA systems. Moreover, we create a benchmark dataset named ICTQA, which consists of QA pairs based on the knowledge of ICT-DATA. This dataset is particularly suitable for evaluating enhanced LLMs, as it includes some industry-specific knowledge not covered in the general LLMs training stage.
@@ 2 Table-to-Text Generation
  • TPLM-based method: This method involves fine-tuning TPLMs on specific task datasets Liu et al. (2022). In this paper, we utilize the MVP model Tang et al. (2023), which initially pre-trains the BART model Lewis et al. (2020a) on numerous natural language generation datasets, followed by fine-tuning on various cross-domain table-to-text datasets. It allows customized adjustment of the output through fine-tuning, offering higher flexibility and domain adaptability, while requiring more computational resources.
@@
Some examples of table-to-text, along with the specific templates and prompts for ChatGPT used in this paper, can be found in Appendix C.
@@ 4.1 Evaluation Dataset
ICT-DATA. We collect ICT-DATA based on 170 English technical documents related to ICT products. Each product document consists of tables and text, whose contents include product descriptions, configuration guides, terms, and definitions, etc. The total storage size is approximately 6GB. Moreover, the text contained in the tables accounts for about 18% of the total dataset content. In Appendix B.2, we provide detailed statistics and the preprocessing methods used for the table data.
ICTQA. We create the ICTQA dataset to evaluate the performance of domain QA systems, by collecting 9,000 questions with long-form answers from the actual ICT product technical support QA platform. All the answers are written by experts based on product documents. We manually select 500 questions as the test set, whose answers involve knowledge from both tables and text. The remaining QA pairs are used as the training set for the instruction fine-tuning phase in the DSFT paradigm. We show statistics and some examples in Appendix B.1.
@@ 4.2 Evaluation Metrics
For fairness and to eliminate potential bias, responses are presented anonymously to both the GPT-4 and human evaluators. The full prompt, evaluation setup for human and scoring criteria are detailed in Appendix E.
@@ 5 Experimental Setup
QA Systems of the DSFT Paradigm. Within the DSFT paradigm, we utilize Meta’s OPT (1.3B to 13B) Zhang et al. (2022) and Llama2-base (7B, 13B) Touvron et al. (2023) as foundation models. The OPT models offer variable sizes to enhance robustness. To mitigate training costs, we employ the QLoRA Dettmers et al. (2023) strategy for pre-training and instruction fine-tuning. The instruction template can be found in Appendix B.3.
@@
Fair Comparison. To maintain consistency and control variables, all models are trained or used under the same settings on four different corpora. Detailed training parameters and GPU costs are available in Appendix D.
@@ 7 Conclusion
Appendix A Related Work
A.1 Domain Augmented Large Language Models.
To enhance the capabilities of LLMs in domain-specific tasks, some research develops LLMs through incremental training on extensive domain-specific corpora Luo et al. (2023); Huang et al. (2023). Other works employ prompt-based solutions, retrieving relevant domain information from external knowledge bases during using LLMs Gao et al. (2023); Wang et al. (2023d); Xu et al. (2023).
A.2 Question Answering over Hybrid Data
Some works study QA tasks on hybrid data that contain both tables and text Zhu et al. (2021); Chen et al. (2020c, a). Popular approaches often involve designing a complex system that has independent modules to process text and tables separately. The information from these two modules is then merged and fed into a language model to generate answers Zhong et al. (2022). Additionally, some methods rely on formulating executable languages (such as SQL) to access tables Nan et al. (2022); Li et al. (2021). These executable languages usually have strict assumptions about the structure of tables, making these methods unsuitable for real-world domain QA system development.
Appendix B ICT Datasets
B.1 ICTQA
To analyze various question types, we follow the classification method of Yang et al. (2023). This approach categorizes questions based on their first interrogative word and assigns tags reflecting the nature of the information sought. The statistical data of this classification is detailed in Table 5. Specifically, the ICTQA dataset labels questions under tags like ‘Parameter’, ‘Configuration’, and ‘Command’, each indicating the type of information requested. For instance, ‘Parameter’ relates to queries about specific values or settings, while ‘Configuration’ pertains to questions regarding the setup of systems or processes. Additionally, questions are grouped by their first interrogative word, with ‘What’, ‘How’, and ‘Why’ being predominant. This categorization sheds light on user inquiries: ‘What’ typically seeks factual details, ‘How’ focuses on procedures or techniques, and ‘Why’ delves into underlying reasons or justifications. The dataset reveals an average question length of 75.13 characters and an average answer length of 160.25 characters. Table 6 shows examples from ICTQA questions.
@@ B.3 Instruction Template
Appendix C Table-to-Text Generation
C.1 Template Design for Table Serialization
@@
C.2 Prompt for LLM-based Method
@@
Table 10: The instruction template.
C.3 Table-to-Text Generation Examples
@@
Table 11: The GPU cost of training each model under the QLoRA strategy using a corpus generated by the markdown table-to-text generation method. The GPU hours are computed as follows: (iteration time (seconds) $\times$ number of iterations $\times$ number of GPUs ÷ 3600 seconds/hour).
Appendix D Training Setup and GPU costs
@@
For the instruction fine-tuning phase, we derive the instruction dataset from the ICTQA training set. The training parameters for this phase are: 5 training epochs, a maximum token length of 512, a batch size of 8 per GPU, and a learning rate of 1e-4 with a cosine decaying scheduler. The configuration of LoRA remains the same as in the pre-training phase.
Appendix E Evaluation Setup
E.1 Prompt for LLM as Evaluator
@@
Table 12 details the prompt template used for evaluating responses via LLM (GPT-4) in a one-shot setting. The task of the GPT-4 is to compare four responses to a grounded answer and assigning each a score ranging from 0 to 5. The score of 0 denoting a punitive score, signifying that the response is marred by repetitive content that diminish linguistic coherence. A score of 1 indicates minimal similarity to the grounded answer, while a score of 5 signifies maximal similarity. Additionally, Table 13 provides an example of the one-shot demonstration used in this evaluation.
E.2 Setup and Criteria for Human Evaluation
@@
For the human evaluation, three co-authors of this paper, all with domain expertise in ICT products, are designated as evaluators. Each sample receives three independent evaluations from these qualified evaluators. We analyze the consistency of scoring across the three evaluations. If the ranking order of the four responses remains consistent and the scoring variance for the same response across different evaluators does not exceed one point, the evaluation is deemed reliable. We compare the scores of the three independent evaluations if the scoring order of the four model responses remains consistent, and the score difference for the same response among different evaluators is no greater than one, we consider this a reliable evaluation. In cases of significant discrepancies, evaluators are requested to reassess the sample. In all evaluation documents, the sources of the responses are anonymously presented as ’A’,’B’,’C’,and ’D’. Table 14 shows the scoring criteria for human evaluation, which is consistent with the scoring criteria for LLM as evaluator presented in Table 12.
@@
Appendix F Case Analysis
We demonstrate a QA case of a RAG QA system built upon the LLaMA-70B-chat model. As shown in Table 18, our RAG QA system successfully retrieves the correct context information containing the query answer from the corpora generated by the Markdown and LLM methods. However, it fails to retrieve correct information from the corpora generated by the TPLM-based and Template methods. We show their t-SNE visualization in the semantic space in Figure 7. In this case, it can be clearly seen that the misleading text chunks generated by the TPLM-based and template method, which are related to entities in the query but do not contain the correct answer, are semantically closer to the query. This leads to the failure to retrieve the correct chunks (i.e., Target chunks in the figure) containing the query’s answer, indicating that the text generated by these two methods has poor semantic representations.
@@ Appendix F Case Analysis
Table 18: A QA example of the RAG QA system based on using a corpus generated by each of the four form-to-text methods as a retrieval source. The red font indicates text in the retrieved passage that is relevant to the answer.

### 09 Apr 2024 v2

@@ 1 Introduction
In terms of implementation, we collect a real-world hybrid dataset called ICT-DATA, by extracting text and tables from numerous documents about Information and Communication Technology (ICT) products. It is important to note that the text contained in tables accounts for approximately 18% of the total content in ICT-DATA (based on word count statistics). We employ different table-to-text methods to process the tables in ICT-DATA, obtaining different ICT corpora. These corpora are then utilized to build QA systems. Moreover, we create a benchmark dataset named ICTQA, which consists of QA pairs based on the knowledge of ICT-DATA. This dataset is particularly suitable for evaluating enhanced LLMs, as it includes some industry-specific knowledge not covered in the general LLMs training stage.
@@ 2 Table-to-Text Generation
  • TPLM-based method: This method involves fine-tuning Traditional Pre-trained Language Models (TPLMs), such as T5 Raffel et al. (2020) and BART Lewis et al. (2020a), on specific table-to-text generation task datasets Liu et al. (2022). In this paper, we utilize the MVP model Tang et al. (2023), which initially pre-trains the BART model on numerous natural language generation datasets, followed by fine-tuning on various cross-domain table-to-text datasets. It allows customized adjustment of the output through fine-tuning, offering higher flexibility and domain adaptability, while requiring more computational resources.
@@
Some examples of table-to-text, along with the specific templates and prompts for ChatGPT used in this paper, can be found in Appendix B.
@@ 4.1 Evaluation Dataset
ICT-DATA. We collect ICT-DATA based on 170 English technical documents related to ICT products. Each product document consists of tables and text, whose contents include product descriptions, configuration guides, terms, and definitions, etc. The total storage size is approximately 6GB. Moreover, the number of words in the table data accounts for about 18% of the total number of words in the dataset. In Appendix A.2, we provide detailed statistics and the preprocessing methods used for the table data.
ICTQA. We create the ICTQA dataset to evaluate the performance of domain QA systems, by collecting 9,000 questions with long-form answers from the actual ICT product technical support QA platform. All the answers are written by experts based on product documents. We manually select 500 questions as the test set, whose answers involve knowledge from both tables and text. The remaining QA pairs are used as the training set for the instruction fine-tuning phase in the DSFT paradigm. We show statistics and some examples in Appendix A.1.
@@ 4.2 Evaluation Metrics
For fairness and to eliminate potential bias, responses are presented anonymously to both the GPT-4 and human evaluators. The full prompt, evaluation setup for human and scoring criteria are detailed in Appendix D.
@@ 5 Experimental Setup
QA Systems of the DSFT Paradigm. Within the DSFT paradigm, we utilize Meta’s OPT (1.3B to 13B) Zhang et al. (2022) and Llama2-base (7B, 13B) Touvron et al. (2023) as foundation models. The OPT models offer variable sizes to enhance robustness. To mitigate training costs, we employ the QLoRA Dettmers et al. (2023) strategy for pre-training and instruction fine-tuning. The instruction template can be found in Appendix A.3.
@@
Fair Comparison. To maintain consistency and control variables, all models are trained or used under the same settings on four different corpora. Detailed training parameters and GPU costs are available in Appendix C.
@@ 6.3 RQ3: Are there practical suggestions for choosing table-to-text methods?
6.4 Additional discussion on experimental results
As shown in Table 2, under the ICT dataset and the experimental setup of this study, the RAG method outperforms the DSFT method in Llama2 models. This demonstrates that RAG has an excellent performance as a lower cost method. We attribute this result to two main reasons: 1). The ICT data used in this study covers dense domain knowledge, and it is still challenging to adapt the LLM well to this complex domain data through incremental pre-training. 2). As the statistical analysis in Appendix A.1, most of the questions in the ICTQA are quizzes on the knowledge of product manuals. In this scenario, the existing excellent dense vector retrievers have high recall accuracy. The studies of Gupta et al. (2024) and Soudani et al. (2024) have respectively conducted detailed experiments on the choice between Fine-Tuning and RAG under the agricultural domain data and Less Popular Knowledge scenarios. Our experimental results in this work further validate their viewpoints. It is also worth noting that in this study, the bge-large-en embedding model Zhang et al. (2023) embeds text chunks into 1024-dimensional vectors. During the retrieval of relevant chunks based on the questions, the peak running memory requirement is approximately 280G.
Another interesting experimental result is that GPT-3.5-turbo performs worse than the Llama2 family in the RAG paradigm. We manually observe the QA cases and find that GPT-3.5-turbo has a significantly higher probability of outputting “I don’t know the answer.”, even if the retriever finds text chunks containing the correct answer.
7 Related Work
7.1 Domain Augmented Large Language Models.
In order to enhance the capabilities of LLMs in domain-specific tasks, some works develop LLMs through incremental training on an extensive domain corpus, inheriting the benefits of both the emergent abilities of LLMs and domain-specific knowledge Luo et al. (2023); Huang et al. (2023). This technology yields significant results, but it demands substantial computational resources and incurs high costs Wang et al. (2023a). In order to overcome this difficulty, a prompt-based solution that does not require updating model parameters has been proposed. They retrieve relevant domain information from external knowledge bases before answering questions with LLMs Gao et al. (2023); Wang et al. (2023d); Xu et al. (2023).
7.2 Question Answering over Hybrid Data
Some works study QA tasks on hybrid data that contain both tables and text Zhu et al. (2021); Chen et al. (2020c, a). Popular approaches often involve designing a complex system that has independent modules to process text and tables separately. The information from these two modules is then merged and fed into a language model to generate answers Zhong et al. (2022). Additionally, some of these methods not only require annotations of metadata identifying text and tables relevant to the question, but they also rely on the formulation of executable languages to access tables, such as SQL or SPARQL Nan et al. (2022); Li et al. (2021). These executable languages often have strict assumptions about the structure of the tables. These limitations make these approaches ill-suited for the real-world LLM-based scenario domain QA systems. Therefore, the results of this study were not compared with these baseline models in the experiments.
@@ 8 Conclusion
Acknowledgements
This work is partially supported by National Nature Science Foundation of China under No. U21A20488. We thank the Big Data Computing Center of Southeast University for providing the facility support on the numerical calculations in this paper.
Appendix A ICT Datasets
A.1 ICTQA
To analyze various question types, we follow the classification method of Yang et al. (2023). This approach categorizes questions based on their first interrogative word and assigns tags reflecting the nature of the information sought. The statistical data of this classification is detailed in Table 5. Specifically, the ICTQA dataset labels questions under tags like ‘Parameter’, ‘Configuration’, and ‘Command’, each indicating the type of information requested. For instance, ‘Parameter’ relates to queries about specific values or settings, while ‘Configuration’ pertains to questions regarding the setup of systems or processes. Additionally, questions are grouped by their first interrogative word. This categorization sheds light on user inquiries: ‘What’ typically seeks factual details, and ‘How’ focuses on procedures or techniques. The average length of questions and answers in ICTQA is 75.13 characters and 160.25 characters, respectively. Table 6 shows examples from ICTQA questions.
@@ A.3 Instruction Template
Appendix B Table-to-Text Generation Setups
B.1 Template Design for Table Serialization
@@
B.2 Prompt for LLM-based Method
@@
Table 10: The prompt template designed for the LLM-based table-to-text generation.
B.3 Table-to-Text Generation Examples
@@
Table 11: The GPU cost of training each model under the QLoRA strategy using a corpus generated by the Markdown table-to-text generation method. The training cost with corpora produced by other methods is close to that of Markdown. The GPU hours are computed as follows: (iteration time (seconds) $\times$ number of iterations $\times$ number of GPUs ÷ 3600 seconds/hour).
Appendix C Training Setup and GPU costs
@@
For the instruction fine-tuning phase, we derive the instruction dataset from the ICTQA training set. The training parameters for this phase are: 5 training epochs, a maximum token length of 512, a batch size of 8 per GPU, and a learning rate of 1e-4 with a cosine decaying scheduler. The configuration of QLoRA remains the same as in the pre-training phase.
Appendix D Evaluation Setup
D.1 Prompt for LLM as Evaluator
@@
Table 12 details the prompt template used for evaluating responses via LLM (GPT-4) in a one-shot setting. The task of the GPT-4 is to compare four responses to a grounded answer and assigning each a score ranging from 0 to 5. Additionally, Table 13 provides an example of the one-shot demonstration used in this evaluation.
D.2 Setup and Criteria for Human Evaluation
@@
For the human evaluation, three co-authors of this paper, all with domain expertise in ICT products, are designated as evaluators. Each sample receives three independent evaluations from these qualified evaluators. We analyze the consistency of scoring across the three evaluations. If the ranking order of the four responses remains consistent and the score difference for the same response across different evaluators does not exceed one point, the evaluation is deemed reliable. In cases of significant discrepancies, evaluators are requested to reassess the sample. In all evaluation documents, the sources of the responses are anonymously presented as ‘A’,‘B’,‘C’, and ‘D’. Table 14 shows the scoring criteria for human evaluation, which is consistent with the scoring criteria for LLM as evaluator presented in Table 12.
@@
Appendix E Case Analysis
We demonstrate a QA case of a RAG QA system built upon the LLaMA-70B-chat model. As shown in Table 18, our RAG QA system successfully retrieves the correct context information containing the query answer from the corpora generated by the Markdown and LLM-based methods. However, it fails to retrieve correct information from the corpora generated by the TPLM-based and Template methods. We show their t-SNE visualization in the semantic space in Figure 7. In this case, it can be clearly seen that the misleading text chunks generated by the TPLM-based and template method, which are related to entities in the query but do not contain the correct answer, are semantically closer to the query. This leads to the failure to retrieve the correct chunks (i.e., Target chunks in the figure) containing the query’s answer, indicating that the text generated by these two methods has poor semantic representations.
@@ Appendix E Case Analysis
Table 18: A QA example of the RAG QA system based on using a corpus generated by each of the four table-to-text methods as a retrieval source. The red font indicates text in the retrieved passage that is relevant to the answer.